OpenStudy (anonymous):

The function h(x) is quadratic and h(3) = h(–10) = 0. Which could represent h(x)? h(x) = x2 – 13x – 30 h(x) = x2 – 7x – 30 h(x) = 2x2 + 26x – 60 h(x) = 2x2 + 14x – 60

OpenStudy (anonymous):

@triciaal

OpenStudy (freckles):

you could plug in 3 and -10 into the h's given to see which would work... or... if you have h(a)=h(b)=0 then x=a and x=b are zeros which means x-a and x-b are factors of h

OpenStudy (anonymous):

@freckles C

OpenStudy (freckles):

can you tell me how you got that

OpenStudy (anonymous):

Pluged In 3 @freckles

OpenStudy (anonymous):

D imean

OpenStudy (freckles):

\[2(3)^2+26(3)-60 \\ 2(9)+78-60 \\ 18+78-60 \neq 0\] ok D makes more since I can also show you the other way I was talking about.. x=3 gives us h is 0 so x-3 is a factor x=-10 gives us h is 0 so x+10 is a factor so we know h(x)=c(x-3)(x+10) where c is not 0 h(x)=c(x^2+7x-30) and our case c is 2 here

OpenStudy (freckles):

sense* (not since :p)

OpenStudy (anonymous):

So D is correct

OpenStudy (triciaal):

h(3) is a zero so x-3 is a factor h(-10) is a zero so x -(-10) is a factor the factors are (x-3) and (x+10) when you multiply these you get the function h(x) = (x-3)(x + 10) remember to watch the signs.

OpenStudy (triciaal):

no the answer would be B

OpenStudy (triciaal):